The total pressure of a mixture of oxygen and hydrogen is 1.00 atm. The mixture is ignited and the water is removed. The remaining gas is pure hydrogen and exerts a pressure of 0.400 atm when measured at the same values of T and V as the original. What is the mole fraction of hydrogen in the original mixture
Answer #1

The original mixture will have 80% of H₂ and 20% of O₂.

Explanation:

Total pressure of the original mixture of oxygen and hydrogen, P1 = 1 atm

Pressure of pure hydrogen, P2= 0.40 atm at same values of T and V as the original mixture

According to the question, we can write

2H₂+ O₂ --> 2H₂O


Let initial number of moles of H₂ is “n1”, Initial number of moles of O₂ be “n2” and initial mole of 2H₂O is zero.

At equilibrium,  

Moles of 2H2 = n1- 2α

Moles of O2 = n2 - α

Moles of 2H2O = 2α  

Since the mixture is ignited therefore the O2 is consumed ∴ n2 – α = 0 <---> n2 = α and for the remaining pure Hydrogen the no. of mole will be n1- 2α = n1 – 2n2.

By ideal gas law, we can write the equations as  

P1 = (n1 + n2)(RT/V) …… (i)

P2 = (n1 – 2n2)(RT/V) …. (ii)


On dividing the eq. (ii) by (i), we get

P2 / P1= [n1/(n1+n2)] – [2n2/(n1+n2​​​​​​)]

[Where, n1/(n1+n2) = X1, mole fraction of hydrogen and n2/(n1+n2) = X2 mole fraction of oxygen]

∴ P2 / P1 = X1– 2(X2) …… (iii)

Or, P2/P1 = 1 – X2 – 2 (X2)  

Or, P2/P1 = 1 – 3 (X2)

Or, 0.40 / 1 = 1 – 3(X2)

Or, 3 (X2) = 1 - 0.40 = 0.60

Or, X2 = 0.60 / 3 = 0.2

Putting value of X₂ in eq. (iii)

∴ 0.4/1 = X₁ – 2 * 0.2

Or, X₁ = 0.4 + 0.4 = 0.8

So X1=0.8 and X2=0.2

∴ Mole fraction of oxygen = 20 % and Mole fraction of hydrogen = 80 %.